# 4Université de Strasbourg, CNRS UMR 7550, Observatoire astronomique de Strasbourg, 67000 Strasbourg, France 5DIO, Observatoire de Paris, CNRS, PSL, 75014, Paris, France 6IDRIS, CNRS, F-91403 Orsay, France

### v1

@@ 3.3 YOLO-CIANNA SDC2 configuration
The typical volume source density in the SDC2 data is very low. Because our output grid already has a high resolution compared to the average source size, we observed that a single unit per grid element is enough to achieve the best score. If trained with more units, the detector often converges toward a solution where a single unit makes all the detections. This is likely due to the small training sample and to the absence of clear and balanced subgroups in the source parameter space (Sect. A). The background detection scaling factor is set to $\lambda_{\rm void}=0.01$.
@@
As stated in Sect. 3.1, our detection units represent objects as bounding boxes, which are essential to the training association function. As thoroughly discussed in Paper I, astronomical sources rarely exhibit the sharp and well-defined edges required to define bounding boxes. We therefore rely on a box proxy of size $S^{\prime}$ in both sky axes and $D$ in the frequency axis for each source (Eq. 11). We rotate this box based on $PA$ and search for the smallest box with edges aligned with the survey axis and that contains all rotated vertices, thereby defining our final target box. The impact of the inclination angle $i$ is already included in the definition of the source HI size. However, the resulting boxes remain small, which increases training difficulty as an unconstrained detector initially relies on fortuitous matches to identify the objects of interest. This problem would arise for every dataset, but it is more pronounced in 3D as positional errors across all dimensions are combined in the DIoU. To compensate, we increase the size of the target bounding boxes by 4 pixels in both sky axes and 10 channels in the frequency axis. We stress that this box definition has no impact on the HI size $S$ and line width prediction $w_{20}$ used in the scorer matching criterion, which are predicted as additional regression parameters. The corresponding target box is illustrated for a few sources in Fig 8. From this, we set our detection unit box size priors to $10\times 10\times 36$. We observed that small changes in the target box sizes can significantly affect the score, whereas changing the size of the prior only has a minimal effect.
@@
Our single detection unit is configured to predict the box center coordinates ($x,y,z$), the box size ($w,h,d$), a detection probability $P$, and an objectness score $O$. To these mandatory elements, we add $f$, $S$, $w_{20}$, $PA$, and $i$, all expressed as individual linearly-activated output vector elements constrained by a regression loss subpart. All parameters are normalized in a similar range to balance their loss contributions. $f$ is clipped in between [10,300] Jy Hz, log-scaled, and rescaled to [0,1]. $S$ is clipped in between [4,25] arcsec and linearly scaled to [0,1]. $w_{20}$ is clipped in between [100,700] km s-1, log-scaled, and also rescaled to [0,1]. To avoid tasking the network with impossible predictions, we set $PA=180^{\circ}$ for all small sources with $S<6$ arcsec (beam size is 7 arcsec). $PA$ is then converted from its [0,360] degree range to two independent parameters in the [-1,1] range through sin and cos functions, which works better than a direct prediction of the angle due to angular symmetries and periodicity. $i$ is converted from its [0,90] range to the [0,1] range through the cos function, which also flattens the otherwise inhomogeneous distribution of $i$. In summary, with six extra parameters, all targets and predictions are expressed in the form of a 14-element vector $\left<x,y,z,w,h,d,P,O,f,S,w_{20},\sin{(PA)},\cos{(PA)},\cos{(i)}\right$>. The individual $\gamma^{p}$ scaling values are given in the top row of Table 2.
@@
The remaining method hyperparameters are listed in Tables 2 and  3. All fIoU thresholds are given for a DIoU-based association function. We note a few differences compared to the SDC1 setup, like the absence of association refinement parameters, which are not required with a single detection unit per grid element. Difficulty-flagging parameters are specified here, but discussed in the next section.
@@
3.3.1 Training example generation and augmentation
@@
Training examples are dynamically generated with augmentation from the normalized LDEV cube (Sect. 2.4) and the selected LDEV truth source catalog (Sect. 2.3). As discussed in Sect. 2.4, we make the approximation that the frequency axis is an invariant spatial coordinate, meaning we train our 3D detector to be insensitive to the input subvolume position on the frequency axis. We define two ways of generating training subcubes. The first one, in 70% of the cases, extracts a cutout at a random position in the LDEV cube regardless of its content. Sources with their central coordinate inside the cutout are added to the target list. This mostly returns empty cubes, but it exposes the network to the full diversity of background signals and artifacts. The second one, in 30% of the cases, draws a random source from the selected truth sample and extracts a random cutout within the subregion of the LDEV cube where the selected source is guaranteed to be included. If other sources have their center in the same region, they are also added to the target list. With this approach, we ensure that a minimum proportion of examples contains at least one detectable source.
@@
Extracted subcubes are augmented to increase the diversity of examples. We add masks on the edges of the subcubes to mimic the effect of added padding when applying the method to the edge of a large cube. Masks are added independently to each axis at a 20% rate, always on a single side picked at random, and masking up to half the subcube in the given axis. We also apply flips at a 50% rate for each axis independently. Whereas this is a standard augmentation for sky axes, it is generally not used in the frequency dimension. It requires assuming that the source morphology is symmetrical over the frequency axis, which is a stronger approximation than treating the source frequency position as invariant. Still, enabling this augmentation results in an improvement in the achievable score of a few percent, indicating that the increased example diversity is more beneficial than the potential drawbacks of this approximation. Target $PA$ values are updated to account for the flips when necessary.
@@
The input resolution of the example subcubes is set to $64\times 64\times 256$. As we discussed in Sect. 3.8 of Paper I, when splitting a finite-sized volume, the choice of chunk size directly impacts the accessible diversity inside each chunk. The smaller the input volume, the higher the diversity, and the lower the chances of overtraining. A small context window also prevents the detector from inferring the position of the input within the large cube, which would cause overtraining. On the other hand, large input sizes are numerically efficient and reduce the proportion of the full volume affected by edge effects and truncated content when making predictions (see Sect. 3.5). We observed that adding augmentations improves the results, indicating that the intrinsic diversity of the LDEV cube is a limiting factor at the selected input size.
@@
Our dynamic augmentation scheme does not fit the classical epoch definition. Instead, we group the generated examples into iterations over 1600 example subcubes. Due to the limited number of labeled sources in the LDEV catalog, splitting them into training and validation sets is impractical. Constituting a large enough validation set so it is statistically relevant would remove too many sources from the training sample. We therefore chose to use all labeled sources from the LDEV catalog for training and relied on occasional scoring over the MAIN SDC2 cube to verify the absence of overtraining, acting as a reliable validation and calibration dataset. During the challenge, the number of scoring was limited to prevent indirect optimization against the hidden MAIN catalog. Yet, it was possible to score a few save states from each trained model to test for blatant overfitting over the LDEV cube. In this paper, we use the scorer more extensively to provide a detailed analysis of our results. Still, it remains prohibitively costly to score more often than every few hundred iterations. Monitoring the training loss directly is not very useful, as most input examples are devoid of visible sources. We instead define a subset composed of cutouts centered on 800 random sources from the training catalog and use it as a proxy validation dataset to monitor the fine-grained loss evolution. Validation loss values are, therefore, only comparable within a given training.
@@ 3.4 Training setup and bootstrapping
Training is performed with the CIANNA framework, which implements our YOLO-CIANNA method, using an RTX 6000 Ada with (91 FP16 TFLOPS). We use FP16C_FP32A mixed-precision training, meaning that computations are done at a 16-bit floating point precision with 32-bit floating point accumulators (Micikevicius et al., 2017). All data flowing through the network uses the reduced 16-bit precision. A master copy of the weight is maintained at 32-bit precision and is used to accumulate the weight updates. Training at this precision yields similar scores to those obtained from full 32-bit training. We set the batch size to 16, which offers a good balance between training efficiency and memory footprint. The learning rate starts at $1.2\times 10^{-5}$ and increases linearly over the first 32 000 examples to $6\times 10^{-4}$. It then decays exponentially as a function of the number of training iterations at a rate of $5\times 10^{-4}$ toward a minimum of $1.2\times 10^{-5}$. The optimizer is a simple mini-batch stochastic gradient descent with a momentum set at 0.7 and a weight decay of $5\times 10^{-4}$. A typical training over 6000 iterations takes around 36 hours on an RTX 6000 Ada, with an average processing speed of $\sim$100 input cubes per second. The best score is generally achieved around iteration 4000.
We observed that small changes in the selection function described in Sect. 2.3 can significantly impact the final result, potentially even preventing training altogether. This is due to an oversimplified selection criterion that does not reflect the real capability of our detector. During training, sources that can be detected but are not included in the target list will be considered false detections, forcing the model to lower the detection probability of all sources that present similar characteristics. Conversely, sources that are undetectable but still included in the target list will force the model to increase the detection probability of background noise. These two types of mislabelling affect the fitting of the objectness curve, resulting in suboptimal detection performances. One possible solution is to rely on the detectors’ self-assessed detection capability to refine the selection function. This is achieved by first training a BASE model from scratch with random weight initialization (using Glorot normal, Glorot & Bengio, 2010) based on our handcrafted selection function. The best iteration is identified using the SDC2 scorer over the MAIN cube following the prediction pipeline. This model is then applied to the LDEV cube that was used for its training to produce a list of detections with a predicted score confidence. We extract the 1600 detections with the highest objectness score. From this, we define a new selection function with lower line-flux and volume brightness thresholds as
@@
but with the additional constraint that the selected sources must match one of the detections produced by the previous model based on a DIoU threshold of 0.1. By doing so, we include fainter true sources considered detectable by the first trained model. However, most added sources are difficult to detect and would likely complicate the training startup of a new model. To compensate, we use the difficult flagging introduced in Appendix A.6.4 of Paper I, and flag all the new sources that do not fulfill the original selection function. This allows the detector to learn from the obvious sources first and progressively refine its solution to include fainter ones. Sources fulfilling the original selection function but not detected by the first model could be removed, but we chose to also flag them as difficult instead. These missed sources might become detectable thanks to the added diversity provided by the new targets. We then train a new bootstrap model based on the refined training source catalog.
This bootstrap process can be repeated until the training sample converges. For the SDC2, we train a BASE model and two successive bootstrap models, BT1 and BT2. Each model is used to refine the following selection function, resulting in training samples containing 1383 (BASE), 1573 (BT1), and 1565 (BT2) target sources. This indicates a form of convergence after a single bootstrap step as the number of selected sources stabilizes, which is confirmed by the best score achieved by each model in Sect. 4.1. Instead of training every new model from scratch, we reuse part of the previously trained model as a starting point. We remove the last three layers from the loaded model and replace them with identically shaped layers with randomly initialized weights. Although this does not improve the final score, it allows faster convergence of bootstrap models.
@@ 3.5 Prediction pipeline
To apply a trained detector to the whole SDC2 MAIN cube, we decompose it into input regions along all three axes. To ensure that a source is fully contained in at least one input region, we need a minimum overlap of at least half the maximum source size in each dimension. In addition, the overlap must be a multiple of the backbone reduction factor to align subregions mapped by detection units from adjacent inputs, which is necessary for multiple detection filtering. We settled on an overlap of 8 pixels in the sky axes and of 32 channels in the frequency axis. This imposes constraints on the detector input size, which must be large enough to minimize the proportion of overlapping volume, as it directly impacts the inference pipeline efficiency. The required wide frequency overlap was one of our motivations in having 256 frequency channels per input (Sects. 3.2 and 3.3.1).
@@
Storing the MAIN cube in system memory to decompose it dynamically is technically challenging. Thus, we rely on the cube slicing described in Sect. 2.4. The detector is applied to all input regions of all subcubes, preserving only detections with an objectness score above 0.1. Multiple detections within each input are filtered through a first non-maximum suppression (NMS) using a 3D DIoU threshold of 0.1. Multiple detections from overlapping inputs are also removed using a secondary NMS with a threshold of -0.3. Detections are then transformed back to the SDC2 catalog format by converting 3D positions to RA, Dec, and central frequency, and by inverting the normalizations for the predicted source parameters.
@@
Classical computer vision tasks are often evaluated in a way that accounts for the detector’s confidence score. In contrast, the SDC2 scorer expects a list of considered real detections. In practice, we can filter our prediction catalog using an objectness threshold deducted from the objectness histogram or by a rough estimate of the expected number of sources. The final threshold can be fine-tuned with a few calls to the scorer to optimize the result, which was our approach during the original challenge. However, to comprehensively analyze our results, we need to perform numerous and comparable scoring over the MAIN cube, which requires an automated threshold search. This is achieved by scoring an unfiltered catalog, from which we extract the list of matches from the scorer products, including the individual score of each source. By binning our catalog by objectness and computing the average score in each bin, we can identify the threshold below which the added sources contribute negatively to the score. Since the detection difficulty likely increases with the frequency, we split the detected sources into 20 frequency bins and optimize independent thresholds for each of them. We acknowledge that this post-process threshold optimization technically breaks the challenge conditions, but it only increases the achievable score by about 1% compared to a manual threshold.
@@
On an RTX 6000 Ada GPU, with inputs of $64\times 64\times 256$, and using the FP16_FP32A compute resolution, we reach a processing speed of 300 inputs per second, or 315 million voxels per second. The GPU memory footprint in inference is around 450 MB per input in the batch, allowing the model to be used for prediction on lighter hardware. The total processing time for the 450 GB datacube is approximately 30 minutes, evenly split between GPU processing and data loading from a high-speed NVMe SSD. This excludes the preprocessing time of the raw MAIN cube, which depends strongly on the available system memory. Post-processing time to filter predictions is negligible.

### v2

@@ 3.3 YOLO-CIANNA SDC2 configuration
The typical volume source density in the SDC2 data is very low. Because our output grid already has a high resolution relative to the average source size, we observed that a single unit per grid element is sufficient to achieve the best score. If trained with more units, the detector often converges toward a solution where a single unit produces all the relevant detections. This is likely due to the small training sample and to the absence of clear and balanced subgroups in the source parameter space (Sect. A). The background detection scaling factor is set to $\lambda_{\rm void}=0.01$.
@@
As stated in Sect. 3.1, our detection units represent objects as bounding boxes, which are essential to the training association function. As thoroughly discussed in Paper I, astronomical sources rarely exhibit the sharp and well-defined edges required to define bounding boxes. We therefore relied on a box proxy of size $S^{\prime}$ in both sky axes and $D$ in the frequency axis for each source (Eq. 11). We rotated this box based on $PA$ and searched for the smallest box whose edges are aligned with the survey axis, and that contains all rotated vertices, thereby defining our final target box. The impact of the inclination angle, $i$, is already included in the definition of the source HI size. However, the resulting boxes remain small, which increases training difficulty as an unconstrained detector initially relies on fortuitous matches to identify the objects of interest. This problem would arise for every dataset, but it is more pronounced in 3D as positional errors across all dimensions are combined in the DIoU. To compensate, we increase the size of the target bounding boxes by 4 pixels in both sky axes and 10 channels in the frequency axis. We stress that this box definition has no impact on the HI size, $S$, and line width prediction, $w_{20}$, used in the scorer matching criterion, which are predicted as additional regression parameters. The corresponding target box is illustrated for a few sources in Fig 8. From this, we set our detection unit box size priors to $10\times 10\times 36$. We observed that small changes in the target box sizes can significantly affect the score, whereas changing the size of the prior has only a minimal effect.
@@
Our single detection unit is configured to predict the box center coordinates ($x,y,z$), the box size ($w,h,d$), a detection probability, $P$, and an objectness score, $O$. To these mandatory elements, we add $f$, $S$, $w_{20}$, $PA$, and $i$, all expressed as individual linearly activated output vector elements constrained by a regression loss subpart. All parameters are normalized in a similar range to balance their loss contributions. $f$ is clipped in between [10,300] Jy Hz, log-scaled, and rescaled to [0,1]. $S$ is clipped in between [4,25] arcsec and linearly scaled to [0,1]. $w_{20}$ is clipped in between [100,700] km s-1, log-scaled, and also rescaled to [0,1]. To avoid tasking the network with impossible predictions, we set $PA=180^{\circ}$ for all small sources with $S<6$ arcsec (beam size is 7 arcsec). $PA$ is then converted from its [0,360] degree range to two independent parameters in the [-1,1] range through sin and cos functions, which works better than a direct prediction of the angle due to angular symmetries and periodicity. $i$ is converted from its [0,90] range to the [0,1] range through the cos function, which also flattens the otherwise inhomogeneous distribution of $i$. In summary, with six extra parameters, all targets and predictions are expressed in the form of a 14-element vector $\left<x,y,z,w,h,d,P,O,f,S,w_{20},\sin{(PA)},\cos{(PA)},\cos{(i)}\right$>. The individual $\gamma^{p}$ scaling values are given in the top row of Table 2.
@@
The remaining method hyperparameters are listed in Tables 2 and 3. All fIoU thresholds are given for a DIoU-based association function. We note a few differences compared to the SDC1 setup, such as the absence of association refinement parameters, which are not required when using a single detection unit per grid element. Difficulty-flagging parameters are specified here, but discussed in the next section.
@@
3.4 Training example generation and augmentation
@@
Training examples were dynamically generated with augmentation from the normalized LDEV cube (Sect. 2.4) and the selected LDEV truth source catalog (Sect. 2.3). As discussed in Sect. 2.4, we made the approximation that the frequency axis is an invariant spatial coordinate, meaning we trained our 3D detector to be insensitive to the input subvolume position on the frequency axis. We defined two ways of generating training subcubes. The first one, in 70% of cases, extracts a cutout at a random position within the LDEV cube regardless of its content. Sources whose central coordinates are within the cutout are added to the target list. This mostly returns empty cubes, but it exposes the network to the full diversity of background signals and artifacts. The second one, in 30% of cases, draws a random source from the selected truth sample and extracts a random cutout within the subregion of the LDEV cube where the selected source is guaranteed to be included. If other sources have their center in the same region, they are also added to the target list. With this approach, we ensure that a minimum proportion of examples contains at least one detectable source.
@@
Extracted subcubes were augmented to increase the diversity of examples. We added masks on the edges of the subcubes to mimic the effect of added padding when applying the method to the edge of a large cube. Masks were added independently to each axis at a 20% rate, always on a single side selected at random, and masking up to half the subcube in that axis. We also applied flips at a 50% rate for each axis independently. Whereas this is a standard augmentation for sky axes, it is generally not used in the frequency dimension. It requires one to assume that the source morphology is symmetric along the frequency axis, which is a stronger approximation than treating the source frequency position as invariant. Still, enabling this augmentation improves the achievable score by a few percent, indicating that the increased example diversity outweighs the potential drawbacks of this approximation. Target $PA$ values are updated to account for flips when necessary.
@@
The input resolution of the example subcubes was set to $64\times 64\times 256$. As we discussed in Sect. 3.8 of Paper I, when splitting a finite-sized volume, the choice of chunk size impacts the accessible diversity inside each chunk. The smaller the input volume, the higher the diversity, and the lower the chances of overtraining. A small context window also prevents the detector from inferring the input’s position within the large cube, which could lead to overtraining. On the other hand, large input sizes are numerically efficient and reduce the proportion of the volume affected by edge effects and truncated content when making predictions (Sect. 3.6). We observed that adding augmentations improves the results, indicating that the intrinsic diversity of the LDEV cube is a limiting factor at the selected input size.
@@
Our dynamic augmentation scheme does not fit the classical definition of an epoch. Instead, we grouped the generated examples into iterations over 1600 example subcubes. Due to the limited number of labeled sources in the LDEV catalog, splitting them into training and validation sets is impractical. Constituting a validation set large enough to be statistically relevant would remove too many sources from the training sample. We therefore chose to use all labeled sources from the LDEV catalog for training and relied on occasional scoring over the MAIN SDC2 cube to verify the absence of overtraining, serving as a validation and calibration dataset. During the challenge, the number of scoring was limited to prevent optimization against the hidden MAIN catalog. Yet, it was possible to score a few saved states for each trained model to test for blatant overfitting over the LDEV cube. In this paper, we used the scorer more extensively to provide a detailed analysis of our results. Still, it remains prohibitively costly to score more often than every few hundred iterations. Monitoring the training loss directly is not very useful, as most input examples lack detectable sources. Instead, we relied on a subset of cutouts centered on 800 randomly selected sources from the training catalog as a proxy validation dataset to monitor fine-grained loss evolution. Validation loss values are therefore only comparable within a given training.
@@ 3.5 Training setup and bootstrapping
Table 4: Training source count over multiple bootstrap steps.
Step Total Match in Match out Similarity
Base 1383 N/A N/A N/A
BT1 1573 1066 190 87.92%
BT2 1565 1090 182 98.60%
BT3 1574 1088 191 99.05%
Training was performed with the CIANNA framework, which implements our YOLO-CIANNA method, using an RTX 6000 Ada with (91 FP16 TFLOPS). We used FP16C_FP32A mixed-precision training, meaning that computations are done at a 16-bit floating point precision with 32-bit floating point accumulators (Micikevicius et al., 2018). All data flowing through the network uses the reduced 16-bit precision. A 32-bit master copy of the weight is maintained and used to accumulate the weight updates. Training at this precision yields similar scores to those obtained from full 32-bit training. We set the batch size to 16, which offers a good balance between training efficiency and memory footprint. The learning rate starts at $1.2\times 10^{-5}$ and increases linearly over the first 32 000 examples to $6\times 10^{-4}$. It then decays exponentially with the number of training iterations at a rate of $5\times 10^{-4}$ toward a minimum of $1.2\times 10^{-5}$. The optimizer is a mini-batch stochastic gradient descent with momentum set at 0.7 and a weight decay of $5\times 10^{-4}$. A typical training over 6000 iterations takes around 36 hours on an RTX 6000 Ada, with an average processing speed of 100 input cubes per second. Best score is generally achieved around iteration 4000.
We observed that small changes to the selection function described in Sect. 2.3 can significantly impact the final result, potentially even preventing training altogether. This is due to an oversimplified selection criterion that does not reflect the detector’s actual capabilities. During training, sources that can be detected but are not included in the target list are considered false detections, forcing the model to lower the detection probability for all sources that present similar characteristics. Conversely, undetectable sources included in the target list force the model to increase the detection probability for background noise. These two types of mislabeling affect the fitting of the objectness curve, resulting in suboptimal detection performances. One possible solution is to rely on the detectors’ self-assessed detection capabilities to refine the selection function. We achieved this by first training a BASE model from scratch with random weight initialization (using Glorot normal, Glorot and Bengio, 2010) based on our handcrafted selection function. The best iteration was identified using the SDC2 scorer over the MAIN cube following the prediction pipeline. This model was then applied to the same LDEV cube used for training in order to produce a list of detections with predicted confidence scores. We extracted the 1600 detections with the highest objectness score. From this, we defined a new selection function with lower line-flux and volume brightness thresholds as
@@
but with the additional constraint that the selected sources must match one of the detections produced by the previous model based on a DIoU threshold of 0.1. By doing so, we included fainter true sources that the first trained model considers detectable. However, most added sources are difficult to detect and would likely complicate the training startup of a new model. To compensate, we used the difficult flagging introduced in Appendix A.6.4 of Paper I and flagged all new sources that do not meet the original selection function. This allows the detector to learn from the obvious sources first and progressively refine its solution to include fainter ones. Sources that fulfill the original selection function but are not detected by the first model could be removed or flagged as difficult. However, we chose to keep them as regular targets, as the score achieved by this first model is already high, indicating no significant negative effect from these sources. Also, these missed sources might become detectable thanks to the added diversity provided by the new targets. The obtained refined target source catalog can then be used to retrain a bootstrap model.
This bootstrap process can be repeated as long as it improves the final score or until the content of the training sample converges. For the SDC2, we trained a BASE model and three successive bootstrap models: BT1, BT2, and BT3. Each model was used to refine the selection function for the next step. Instead of training every new model from scratch, we reused part of the previously trained model as a starting point. The last three layers of the loaded model were replaced with identically shaped layers but with randomly initialized weights. We show and discuss what happens when bootstrapping with a from-scratch model in Appendix D. In table 4, we indicate the evolution of the number of training sources for each bootstrap step, the number of sources from the original selection function for which a match is found by the previous model, the number of fainter sources from the new selection function for which a match is found that are added to the training sample, and the fraction of the final sources that were already present in the training sample of the previous step. From this, we can see diminishing returns as we add bootstrap steps. The first step yields the greatest change in the training sample content, whereas there is only a 1% difference (about ten sources) between BT2 and BT3 training samples. Overall, the training content is so similar across all steps after the first one that it is unlikely to result in significant score differences. This is mostly confirmed by the fact that the remaining score difference between our bootstrap models presented in Sect. 4.1 is of the same order as the retraining variability on a given training dataset. The exact best-scoring step after the first one is mostly random. While only a few steps are sufficient in this specific context, the optimal number of bootstrap steps might be much larger on other datasets and strongly depend on the initial selection function and the exact strategy used to reinject new targets. These elements are further discussed in Appendix E, where we evaluate the effect of starting the training from a degraded initial selection function.
@@ 3.6 Prediction pipeline
Figure 4: Evolution of the validation loss subparts (natural, Sect. A.7 in Paper I) during training for our four successive bootstrap models. The first 20 iterations are omitted. Scores computed over the MAIN cube are indicated every 200 iterations after 400.
@@
To apply a trained detector to the whole SDC2 MAIN cube, we decomposed it into input regions along all three axes. To ensure that a source is fully contained in at least one input region, a minimum overlap of at least half the maximum source size is required in each dimension. In addition, the overlap must be a multiple of the backbone reduction factor to align subregions mapped by detection units from adjacent inputs, which is necessary for multiple detection filtering. We settled on an overlap of 8 pixels along the sky axes and of 32 channels along the frequency axis. This imposes constraints on the detector input size, which must be large enough to minimize the proportion of overlapping volume, as this directly affects the computational efficiency of the inference pipeline. The required wide frequency overlap was one of the motivations for having 256 frequency channels per input (Sects. 3.2 and 3.4).
@@
Storing the MAIN cube in system memory and decomposing it dynamically is technically challenging. Thus, we rely on the cube slicing described in Sect. 2.4. The detector is applied to all input regions of all subcubes, preserving only detections with an objectness score above 0.1. Multiple detections within each input are filtered through a first non-maximum suppression (NMS) using a 3D DIoU threshold of 0.1. Multiple detections from overlapping inputs are also removed using a secondary NMS with a threshold of -0.3. Detections are then transformed back to the SDC2 catalog format by converting 3D positions to RA, Dec, and central frequency, and by inverting the normalization applied to the predicted source parameters.
@@
Classical computer vision tasks are often evaluated in a way that accounts for the detector’s confidence score. In contrast, the SDC2 scorer expects a list of considered real detections. In practice, we can filter our prediction catalog using an objectness threshold deducted from the objectness histogram or by a rough estimate of the expected number of sources. The final threshold can be fine-tuned with a few calls to the scorer to optimize the result, which was our approach during the original challenge. However, to comprehensively analyze our results, we need to perform numerous comparable scoring on the MAIN cube, which requires an automated threshold search. This is achieved by scoring an unfiltered catalog, from which we extract the list of matches from the scorer products, including the individual score of each source. By binning our catalog by objectness and computing the average score in each bin, we can identify the threshold below which adding sources decreases the score. Since detection difficulty likely increases with frequency, we split the detected sources into 20 frequency bins and optimize independent thresholds for each bin. We acknowledge that this post-process threshold optimization technically breaks the challenge conditions, yielding a near 1% improvement in maximum achievable score over a naive threshold selection. Still, this difference remains much smaller than the typical score difference observed between independent detection methods and allows us to compare all models produced with YOLO-CIANNA in their optimal selection regime, ensuring that the score difference between different setups or bootstrap steps are significative.
@@
On an RTX 6000 Ada GPU, with inputs of $64\times 64\times 256$ and using FP16 compute resolution, we achieve a processing speed of 300 inputs per second, or 315 million voxels per second. The GPU memory footprint during inference is around 450 MB for a batch size of one and scales mostly linearly with the batch size, allowing the model to be used for prediction on entry-level hardware. The total processing time for the 450 GB datacube is approximately 30 minutes, evenly split between GPU processing and data loading from a high-speed NVMe SSD. This excludes the preprocessing time of the raw MAIN cube, which depends strongly on the available system memory. Post-processing time to filter the model’s predictions is negligible.